OpenStudy (anonymous):

X^2+4x x^2=-3x-4 x^2-2x+7=7+x

OpenStudy (anonymous):

can anyone show me how to solve these quadratics

OpenStudy (joshawad):

not pre-algebra sorry

OpenStudy (anonymous):

What is the first one? The second equation can be solved by following: \[x^2 + 3x + 4 = 0\] \[\Delta = 3^2 - 4*4 = -7 < 0 \rightarrow \] No solution (or imaginary solution) The third one is: \[x^2 - 3x = 0\] then \[x_1 = 0\] or \[x_2 = 3\] That's all

OpenStudy (anonymous):